OpenStudy (anonymous):

simplify 4z^2 - 16z + 15 -------------- 2z^2 -11z +15

sam (.sam.):

\[\frac{4z^2 - 16z + 15}{2z^2 -11z +15}\] Factor, \[\frac{(2z-3)(2z-5)}{(2z-5)(z-3)}\] (2z-5) cancels, Leaving \[\frac{2z-3}{z-3}\]

OpenStudy (anonymous):

oh wow thanks!